# Positivstellensätze for polynomial matrices with universal quantifiers

### v2

@@ 7. Conclusions
We extend various classical Positivstellensätze to provide SOS-structured characterizations for polynomial matrices that are positive (semi)definite over a semialgebraic set defined by a PMI with UQs. Under the Archimedean condition, we first present a Scherer-Hol-type Positivstellensatz, along with a sparse version for scalar polynomials being positive over the semialgebraic set in the presence of correlative sparsity. We also derive Putinar–Vasilescu-type, Pólya-type, and Lasserre-Netzer-type Positivstellensätze without assuming the Archimedean condition. These results significantly extend existing work on Positivstellensätze, and some of them remain novel and intriguing even in the absence of UQs. Our results are valuable for ensuring robustness of PMIs over a prescribed set with uncertainty, allowing potential applications in areas such as optimal control, systems theory, and certifying stability.
@@ acknowledgements
Feng Guo was supported by the Chinese National Natural Science Foundation under grant 12471478. Jie Wang was supported by National Key R&D Program of China under grant No. 2023YFA1009401, the Strategic Priority Research Program of the Chinese Academy of Sciences XDB0640000 & XDB0640200, and the National Natural Science Foundation of China under grant No. 12201618 & 12171324.

### v4

@@ 7. Conclusions
We extend various classical Positivstellensätze to provide SOS-structured characterizations for polynomial matrices that are positive (semi)definite over a semialgebraic set defined by a PMI with UQs. Under the Archimedean condition, we first present a matrix-valued Positivstellensatz incorporating universal quantifiers, along with a sparse version for scalar-valued polynomial objectives, leveraging the correlative sparsity patterns. We also derive two generalized Positivstellensätze without assuming the Archimedean condition. These results significantly extend the existing work on Positivstellensätze, and some of them remain novel and intriguing even in the absence of UQs. Our results are valuable in ensuring the robustness of PMIs over a prescribed set with uncertainty, allowing potential applications in areas such as optimal control, systems theory, and certifying stability.
@@ Acknowledgements
The authors appreciate Professor Konrad Schmüdgen for providing the result in [44, Theorem 10.25], which significantly simplified the proof of Theorem 3.1. Feng Guo was supported by the Chinese National Natural Science Foundation under grant 12471478. Jie Wang was supported by National Key R&D Program of China under grant No. 2023YFA1009401, the Strategic Priority Research Program of the Chinese Academy of Sciences XDB0640000 & XDB0640200, and the National Natural Science Foundation of China under grant No. 12201618 & 12171324.